# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 5.1.2. Independent of the choice of the polarization
Let $X$ be a compact Kähler klt space of dimension $n$, $\pi:\widehat{X}\rightarrow X$ be a resolution of singularities of $X$ and $(\mathcal{E}_{X},\theta_{X})$ be a locally free Higgs sheaf over $X$. Denote $(\mathcal{E}_{\widehat{X}},\theta_{\widehat{X}})=\pi^{*}(\mathcal{E}_{X},\theta_{X})$. Then the following are equivalent:
@@
    There exists some Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ such that $(\mathcal{E}_{X},\theta_{X})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable with (1.3).
@@
    ${\mathrm{ch}}_{1}(\mathcal{E}_{\widehat{X}})\cdot[\alpha_{0}]\cdots[\alpha_{n-2}]=0\ \text{and}\ {\mathrm{ch}}_{2}(\mathcal{E}_{\widehat{X}})=0.$
@@
    All Chern classes of $\mathcal{E}_{X}$ vanish and $(\mathcal{E}_{X},\theta_{X})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable with respect to any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$.
Ditto for polystability.
On the other hand, using Theorem 5.5 together with Theorem 5.4, Lemmas 2.21 and 2.24, one can easily verify that the definitions of the categories $\mathrm{Higgs}_{X},\mathrm{pHiggs}_{X},\mathrm{Higgs}_{X_{\mathrm{reg}}}$ and $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ are independent of the choice of the polarization $(\omega_{0},\cdots,\omega_{n-2})$.
@@
The implication $(iv)\Rightarrow(i)$ is obvious and $(iii)\Rightarrow(iv)$ follows from the item (4.1.1). Let us first verify $(i)\Rightarrow(ii)$. Let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover of $X$, $\widehat{Y}$ be a desingularisation of the (unique) irreducible component of the fibre product $\widehat{X}\times_{X}Y$ and then we have the following commuting diagram
@@
(5.5) $\begin{split}&\left(\mathrm{ch}_{2}(\mathcal{Q}_{k}^{\vee\vee})-\frac{1}{\operatorname{rank}\mathcal{Q}_{k}^{\vee\vee}}\cdot c_{1}(\mathcal{Q}_{k}^{\vee\vee})^{2}\right)\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\end{split}$
by letting $\delta\rightarrow 0$. Combining Proposition 2.15, the induction hypothesis (5.5) and the Hodge index Theorem (Lemma 5.3), it holds that
(5.6) $\begin{split}0&=\mathrm{ch}_{2}(\mathcal{E}_{X})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\\ &\leq\sum\limits_{k=1}^{l}2\cdot\mathrm{ch}_{2}(\mathcal{Q}_{k}^{\vee\vee})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\\ &\leq\sum\limits_{k=1}^{l}\frac{1}{\operatorname{rank}\mathcal{Q}_{k}^{\vee\vee}}\cdot c_{1}(\mathcal{Q}_{k}^{\vee\vee})^{2}\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\\ &\leq\sum\limits_{k=1}^{l}\frac{\big(c_{1}(\mathcal{Q}_{k}^{\vee\vee})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m}]\cdots[\alpha_{n-2}]\big)^{2}}{\operatorname{rank}\mathcal{Q}_{k}^{\vee\vee}\cdot[\pi^{*}\alpha_{m}]^{2}\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]}=0.\end{split}$
@@
The proof of the polystable case is similar. ∎
Combining Theorem 5.5 with the classical non-abelian Hodge correspondence for a single Kähler class [17, 97, 86, 23] (see also [93, Theorem 1.9] for projective flat bundles), we immediately obtain its extension to multiple Kähler polarization on compact Kähler manifolds, which has already been established in the polystable case by [11].
@@
This equivalence of categories in this theorem is indeed functorial with respect to the morphisms between compact Kähler manifolds, which can be easily seen from the construction of $\mu_{X}$ (see e.g. [93, Section 9]).
@@
Now let us show Theorem 1.15. Suppose that $f:Z\rightarrow X$ is a bimeromorphic, holomorphic map between compact Kähler klt spaces and $(\mathcal{E}_{Z},\theta_{Z})\in\mathrm{Higgs}_{Z}$, we need to construct a $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$ such that $\pi^{*}(\mathcal{E}_{X},\theta_{X})=(\mathcal{E}_{Z},\theta_{Z})$.
Let $g:W\rightarrow X$ be an orbifold modification of $X$ and $Y$ be a desingularisation of the (unique) irreducible component of the fibre product $W\times_{X}Z$. Then we have the following commuting diagram
@@
In particular, $\pi_{1}:Y\rightarrow W$, $\pi_{2}:Y\rightarrow Z$ and $\pi:Y\rightarrow X$ are resolution of singularities of $Z$, $W$ and $X$, respectively. Recall $(\mathcal{E}_{Y},\theta_{Y}):=\pi_{2}^{*}(\mathcal{E}_{Z},\theta_{Z})\in\mathrm{Higgs}_{Y}$ by Theorem 5.5. It reduces to show the following statement by Lemma (4.1.1).
@@
There exists some locally free Higgs sheaf $(\mathcal{E}_{X},\theta_{X})$ on $X$ such that $(\mathcal{E}_{Y},\theta_{Y})=\pi^{*}(\mathcal{E}_{X},\theta_{X})$
Let us set the stage with some preparations. Suppose that $E_{Y}=\mu_{Y}(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{LSys}_{Y}$ and define $E_{X}=(\pi^{*})^{-1}E_{Y}\in\mathrm{LSys}_{X}$. Denote by $E_{\rho_{X}}$ and $E_{\rho_{Y}}$ the corresponding local systems.
Suppose that that $E_{X}$ is induced by a linear representation $\rho_{X}:\pi_{1}(X)\rightarrow\mathrm{GL}(r,\mathbb{C})$ in block upper-triangular form
@@
$0=E_{0}\subsetneq E_{1}\subsetneq\cdots\subsetneq E_{m}=E_{X},$
where each $E_{i}$ corresponds to the truncated representation
@@
(5.8) $0\subsetneq E_{W,1}\subsetneq\cdots\subsetneq E_{W,m}=E_{W}$
such that each $E_{W,l}=g^{*}E_{X,l}$ is induced by $\widehat{\rho_{i}}\circ g_{*}$. Because $g_{*}:\pi_{1}(W)\rightarrow\pi_{1}(X)$ is an isomorphism by Takayama’s result [101], the quotient
@@
is a simple local system induced by $\rho_{i}\circ g_{*}$. Analogously, the flat bundle $(V_{Y},D_{Y})$ induced by $E_{Y}$ admits a filtration of flat bundles
(5.9) $0\subsetneq V_{Y,1}\subsetneq\cdots\subsetneq V_{Y,m}=V_{Y}$
such that each $V_{Y,l}$ is $D$-flat and induced by $\widehat{\rho_{i}}\circ\pi_{*}$, and the quotient $Q_{Y,i}=V_{Y,i}/V_{Y,i-1}$ with the induced flat connection $D_{Y,i}$ is induced by $\pi^{*}\rho_{i}\in\mathrm{sRep}_{Y}$.
By applying non-abelian Hodge correspondence for compact Kähler manifolds in [23], $(Q_{Y,i},\theta_{Y,i})$ admits a harmonic metric $h_{i}$. Recall there is a unique harmonic decomposition
$D_{Y,i}=\partial_{Y,i}+\bar{\partial}_{Y,i}+\theta_{Y,i}+\theta_{Y,i}^{*}$
and $Q_{Y_{i}}$ can be associate with a Higgs bundle structure $(\mathcal{Q}_{Y,i},\theta_{Y,i})=(Q_{Y,i},\bar{\partial}_{Y,i},\theta_{Y,i}).$ Building on Proposition 4.11 and Theorem 1.12, there exists a locally Higgs sheaf $(\mathcal{Q}_{X,i},\theta_{X,i})\in\mathrm{pHiggs}_{X}$ on $X$ such that $(\mathcal{Q}_{Y,i},\theta_{Y,i})=\pi^{*}(\mathcal{Q}_{X,i},\theta_{X,i}).$ According to the above construction, the underlying $C^{\infty}$-bundle of $(\mathcal{Q}_{X,i},\theta_{X,i})$ is isomorphic to the underlying $C^{\infty}$-bundle of $Q_{X,i}$. In particular,
$(\mathcal{Q}_{W,i},\theta_{W,i}):=g^{*}(\mathcal{Q}_{X,i},\theta_{X,i})\in\mathrm{pHiggs}_{W}$
and its underlying $C^{\infty}$-bundle is isomorphic to the underlying $C^{\infty}$-bundle of $Q_{W,i}$ ($C^{\infty}$ makes sense via local embeddings of analytic spaces).
Now let us complete the proof of Proposition 5.8. The basic idea is constructing a locally free Higgs sheaf $(\mathcal{E}_{W},\theta_{W})$ of the underlying $C^{\infty}$-bundle of $E_{\rho_{W}}$ via the filtration $\{E_{W,k}\}_{k=1}^{m}$ over a standard orbifold structure $W_{\mathrm{orb}}=\{U_{\alpha},G_{\alpha},\mu_{\alpha}\}$ of $W$ and showing $(g_{*}\mathcal{E}_{W})^{\vee\vee}$ is locally free.
@@
Let $(V_{\mathrm{orb}},D_{\mathrm{orb}})$ be the flat orbi-bundle induced by $\{\mu_{\alpha}^{*}E_{\rho_{W}}\}$. It admits a filtration of flat orbi-bundles
@@
such that each $V_{\mathrm{orb},i}$ in invariant under $D_{\mathrm{orb}}$ and the quotient $Q_{\mathrm{orb},i}=V_{\mathrm{orb},i}/V_{\mathrm{orb},i}$ with the induced flat connection $D_{\mathrm{orb},i}$ is induced by $\{\mu_{\alpha}^{*}Q_{W_{i}}\}$. Note that $Q_{\mathrm{orb},i}$ has a Higgs orbi-bundle structure $\{Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i}\}$ induced by $\{\mu_{\alpha}^{*}(\mathcal{Q}_{W,i},\theta_{W,i})\}$. Then $\{Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i}\}$ is polystable with respect to some orbifold Kähler class $\omega_{\mathrm{orb}}$ by Theorem 5.5 and [110, Proposition 4.8], and all orbifold Chern classes vanish because $Q_{\mathrm{orb},i}$ is flat. Building on the existence of the Hermitian-Einstein metric of stable Higgs orbi-bundles due to [110], $(Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i})$ admits an orbifold pluri-harmonic metric $H_{\mathrm{orb},i}$. Since $\mu_{\alpha}$ is étale over $W_{\mathrm{reg}}$, then $H_{\mathrm{orb},i}$ descend to a pluri-harmonic metric $H_{i}$ of $(\mathcal{Q}_{W,i},\theta_{W,i})|_{W_{\mathrm{reg}}}$. By virtue of the uniqueness by Theorem 1.13 (1), $D_{\theta_{W,i},H_{i}}$ identifies with the flat connection structure of $Q_{W,i}|_{W_{\mathrm{reg}}}$ and so $D_{\theta_{\mathrm{orb},i},H_{\mathrm{orb},i}}=D_{\mathrm{orb},i}.$
Note that the Hodge isomorphism Theorem holds in the context of compact complex orbifolds (cf. [77, Proposition 2.1]), so we could completely follow the argument of [23, Proof of Theorem A] to construct a Higgs orbi-bundle structure $(V_{\mathrm{orb}},\bar{\partial}_{\mathrm{orb}},\theta_{\mathrm{orb}})$ of $V_{\mathrm{orb}}$ such that each $V_{\mathrm{orb},i}$ is $\bar{\partial}_{\mathrm{orb}}$-invariant and $\theta_{\mathrm{orb}}$-invariant, and the induced Higgs bundle structures on $Q_{\mathrm{orb},i}$ coincides with $(Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i})$. Consequently,
@@
is the extension of locally Higgs sheaf $\{\mu_{\alpha}^{*}(\mathcal{Q}_{W,i},\theta_{W,i})\}_{i=1}^{m}$. By applying [33, Lemma A.4], $\left((\mu_{\alpha})_{*}(\mu_{\alpha}^{*}\mathcal{Q}_{W,i})\right)^{G_{\alpha}}$ is reflexive and coincides with the locally free sheaf $\mathcal{Q}_{W,i}$ since $\mu_{\alpha}$ is étale over $W_{\mathrm{reg}}$; $\mathcal{E}_{W}:=\left((\mu_{\alpha})_{*}\mathcal{E}_{\alpha}\right)^{G_{\alpha}}$ is reflexive and admits a filtration of saturated subsheaves

### v2

@@ 5.1.2. Independence of the choice of Kähler polarization
Let $X$ be a compact Kähler klt space of dimension $n$, $\pi:\widehat{X}\rightarrow X$ a resolution of singularities, and $(\mathcal{E}_{X},\theta_{X})$ a locally free Higgs sheaf over $X$. Set $(\mathcal{E}_{\widehat{X}},\theta_{\widehat{X}})=\pi^{*}(\mathcal{E}_{X},\theta_{X})$. Then the following statements are equivalent:
@@
    There exist Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ such that $(\mathcal{E}_{X},\theta_{X})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable and satisfies (1.3).
@@
    ${\mathrm{ch}}_{1}(\mathcal{E}_{\widehat{X}})\cdot[\alpha_{0}]\cdots[\alpha_{n-2}]=0\quad\text{and}\quad{\mathrm{ch}}_{2}(\mathcal{E}_{\widehat{X}})=0.$
@@
    All Chern classes of $\mathcal{E}_{X}$ vanish, and $(\mathcal{E}_{X},\theta_{X})$ is semistable with respect to every Kähler polarization $(\omega_{0},\cdots,\omega_{n-2})$.
The same equivalences hold for polystability.
On the other hand, using Theorem 5.5 together with Theorem 5.4, Proposition 2.21 and Lemma 2.24, one can easily verify that the definitions of the categories $\mathrm{Higgs}_{X},\mathrm{pHiggs}_{X},\mathrm{Higgs}_{X_{\mathrm{reg}}}$ and $\mathrm{pHiggs}_{X_{\mathrm{reg}}}$ are independent of the choice of Kähler polarization $(\omega_{0},\cdots,\omega_{n-2})$.
@@
The implication $(iv)\Rightarrow(i)$ is obvious, and $(iii)\Rightarrow(iv)$ follows from item (4.1.a). We first prove $(i)\Rightarrow(ii)$. Let $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover of $X$, $\widehat{Y}$ be a desingularization of the (unique) irreducible component of the fibre product $\widehat{X}\times_{X}Y$. Then we have a commutative diagram:
@@
(5.5) $\begin{split}&\left(\mathrm{ch}_{2}(\mathcal{Q}_{k}^{\vee\vee})-\frac{1}{\operatorname{rank}\mathcal{Q}_{k}^{\vee\vee}}\cdot c_{1}^{2}(\mathcal{Q}_{k}^{\vee\vee})\right)\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\leq 0.\end{split}$
Combining Proposition 2.15, the induction hypothesis for m1m-1, (5.5) and the Hodge Index Theorem (Lemma 5.3), we obtain
(5.6) $\begin{split}0&=\mathrm{ch}_{2}(\mathcal{E}_{\widehat{X}})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\\ &\leq\sum\limits_{k=1}^{l}2\cdot\mathrm{ch}_{2}(\mathcal{Q}_{k}^{\vee\vee})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\\ &\leq\sum\limits_{k=1}^{l}\frac{1}{\operatorname{rank}\mathcal{Q}_{k}^{\vee\vee}}\cdot c_{1}^{2}(\mathcal{Q}_{k}^{\vee\vee})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]\\ &\leq\sum\limits_{k=1}^{l}\frac{\big(c_{1}(\mathcal{Q}_{k}^{\vee\vee})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m}]\cdots[\alpha_{n-2}]\big)^{2}}{\operatorname{rank}\mathcal{Q}_{k}^{\vee\vee}\cdot[\alpha_{m}]^{2}\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]}=0.\end{split}$
Thus, inequality in the fourth line of (5.6) is an equality, which forces that for all $1\leq k\leq l$,
(5.7) $c_{1}^{2}(\mathcal{Q}_{k}^{\vee\vee})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]=\frac{\big(c_{1}(\mathcal{Q}_{k}^{\vee\vee})\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m}]\cdots[\alpha_{n-2}]\big)^{2}}{[\alpha_{m}]^{2}\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]}.$
Combining this with Lemma 5.3, we deduce that for every $1\leq k\leq l$,
(5.8) $c_{1}(\mathcal{Q}_{k}^{\vee\vee})\cdot[\alpha_{m,\epsilon}]\cdot[\alpha_{0,\epsilon}]\cdots[\alpha_{m-1,\epsilon}]\cdot[\alpha_{m+1}]\cdots[\alpha_{n-2}]=0.$
@@
The proof for the polystable case is identical. ∎
Combining Theorem 5.5 with the classical non-abelian Hodge correspondence for a single Kähler class [17, 100, 88, 23] (see also [95, Theorem 1.9] for projective flat bundles), we immediately obtain its extension to multiple Kähler polarizations on compact Kähler manifolds, which has already been established in the polystable case by [11].
@@
The equivalence in Theorem 5.6 is indeed functorial with respect to morphisms between compact Kähler manifolds, which can be seen directly from the construction of $\mu_{X}$ (see e.g. [95, Section 9]).
@@
We now prove Theorem 1.15. Suppose that $f:Z\rightarrow X$ is a bimeromorphic holomorphic map between compact Kähler klt spaces and $(\mathcal{E}_{Z},\theta_{Z})\in\mathrm{Higgs}_{Z}$. We need to construct a locally free Higgs sheaf $(\mathcal{E}_{X},\theta_{X})\in\mathrm{Higgs}_{X}$ such that $f^{*}(\mathcal{E}_{X},\theta_{X})=(\mathcal{E}_{Z},\theta_{Z})$.
Let $g:W\rightarrow X$ be an orbifold modification of $X$ and $Y$ be a desingularization of the (unique) irreducible component of the fibre product $W\times_{X}Z$. Then we have the following commutative diagram
@@
where $pr_{1},pr_{2}$ are morphisms induced by natural projections and $\pi:=g\circ pr_{1}=f\circ pr_{2}$. In particular, $pr_{1}:Y\rightarrow W$, $pr_{2}:Y\rightarrow Z$ and $\pi:Y\rightarrow X$ are resolutions of singularities of $W$, $Z$ and $X$, respectively. By Theorem 5.5, the pullback $(\mathcal{E}_{Y},\theta_{Y}):=pr_{2}^{*}(\mathcal{E}_{Z},\theta_{Z})\in\mathrm{Higgs}_{Y}$. Combining this with item (4.1.a), it reduces to show the following statement.
@@
There exists a locally free Higgs sheaf $(\mathcal{E}_{X},\theta_{X})$ on $X$ such that $(\mathcal{E}_{Y},\theta_{Y})=\pi^{*}(\mathcal{E}_{X},\theta_{X})$.
We begin with some preparations. Assume that $E_{Y}=\mu_{Y}(\mathcal{E}_{Y},\theta_{Y})\in\mathrm{LSys}_{Y}$. Then by item (4.1.b), $E_{X}=(\pi^{*})^{-1}E_{Y}\in\mathrm{LSys}_{X}$ is well-defined. Denote by $E_{\rho_{X}}$ and $E_{\rho_{Y}}$ the corresponding local systems.
Suppose that $E_{X}$ is induced by a linear representation $\rho_{X}:\pi_{1}(X)\rightarrow\mathrm{GL}(r,\mathbb{C})$ in block upper-triangular form
@@
$0=E_{X,0}\subsetneq E_{X,1}\subsetneq\cdots\subsetneq E_{X,m}=E_{X},$
where each $E_{X,i}$ corresponds to the truncated representation
@@
(5.10) $0=E_{W,0}\subsetneq E_{W,1}\subsetneq\cdots\subsetneq E_{W,m}=E_{W}$
such that each $E_{W,i}=g^{*}E_{X,i}$ is induced by $\widehat{\rho_{i}}\circ g_{*}$. Because $g_{*}:\pi_{1}(W)\rightarrow\pi_{1}(X)$ is an isomorphism by item (4.1.b), the quotient
@@
is a simple local system induced by $\rho_{i}\circ g_{*}$. Analogously, the flat bundle $(V_{Y},D_{Y})$ induced by $E_{Y}$ admits a filtration by flat subbundles
(5.11) $0=V_{Y,0}\subsetneq V_{Y,1}\subsetneq\cdots\subsetneq V_{Y,m}=V_{Y}$
such that each $V_{Y,i}$ is $D_{Y}$-flat and induced by $\widehat{\rho_{i}}\circ\pi_{*}$, and the quotient $Q_{Y,i}=V_{Y,i}/V_{Y,i-1}$ with the induced flat connection $D_{Y,i}$ is induced by $\pi^{*}\rho_{i}\in\mathrm{sLSys}_{Y}$.
By applying the non-abelian Hodge correspondence for compact Kähler manifolds [23], $(Q_{Y,i},D_{Y,i})$ admits a pluri-harmonic metric $h_{i}$. Recall that there is a unique harmonic decomposition
(5.12) $D_{Y,i}=\partial_{Y,i}+\bar{\partial}_{Y,i}+\theta_{Y,i}+\theta_{Y,i}^{*},$
and $Q_{Y,i}$ thereby inherits a natural Higgs bundle structure $(\mathcal{Q}_{Y,i},\theta_{Y,i})$. By Proposition 4.11 and Theorem 1.12, there exists a locally free Higgs sheaf $(\mathcal{Q}_{X,i},\theta_{X,i})\in\mathrm{pHiggs}_{X}$ on $X$ such that $(\mathcal{Q}_{Y,i},\theta_{Y,i})=\pi^{*}(\mathcal{Q}_{X,i},\theta_{X,i}).$ According to the above construction, the underlying $C^{\infty}$-bundle of $(\mathcal{Q}_{X,i},\theta_{X,i})$ is isomorphic to the underlying $C^{\infty}$-bundle of $Q_{X,i}$. In particular,
$(\mathcal{Q}_{W,i},\theta_{W,i}):=g^{*}(\mathcal{Q}_{X,i},\theta_{X,i})\in\mathrm{pHiggs}_{W},$
and its underlying $C^{\infty}$-bundle is isomorphic to the underlying $C^{\infty}$-bundle of $Q_{W,i}$ (the notion of a $C^{\infty}$-bundle makes sense via local embeddings of analytic spaces).
Now we complete the proof of Proposition 5.8. The basic idea is to construct a locally free Higgs sheaf structure $(\mathcal{E}_{W},\theta_{W})$ on the underlying $C^{\infty}$-bundle of $E_{\rho_{W}}$ via the filtration $\{E_{W,k}\}_{k=1}^{m}$ and the standard orbifold structure $W_{\mathrm{orb}}=\{(U_{\alpha},G_{\alpha},\mu_{\alpha})\}$ of $W$, and then show that $(g_{*}\mathcal{E}_{W})^{\vee\vee}$ is locally free.
@@
Let $(V_{\mathrm{orb}},D_{\mathrm{orb}})$ be the flat orbi-bundle induced by $\{\mu_{\alpha}^{*}E_{\rho_{W}}\}$. Because $\mu_{\alpha}$ is quasi-étale, $(V_{\mathrm{orb}},D_{\mathrm{orb}})$ admits a filtration by flat orbi-bundles
@@
such that each $V_{\mathrm{orb},i}$ is invariant under $D_{\mathrm{orb}}$ and the quotient $Q_{\mathrm{orb},i}=V_{\mathrm{orb},i}/V_{\mathrm{orb},i-1}$ with the induced flat connection $D_{\mathrm{orb},i}$ is induced by $\{\mu_{\alpha}^{*}Q_{W,i}\}$. Note that $Q_{\mathrm{orb},i}$ carries a Higgs orbi-bundle structure $\{(Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i})\}$ induced by $\{\mu_{\alpha}^{*}(\mathcal{Q}_{W,i},\theta_{W,i})\}$. Then $(Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i})$ is polystable with respect to some orbifold Kähler class $\omega_{\mathrm{orb}}$ by Theorem 5.5 and [113, Proposition 4.8], and all its orbifold Chern classes vanish because $Q_{\mathrm{orb},i}$ is flat. By the existence of Hermitian-Einstein metrics on stable Higgs orbi-bundles [113], $(Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i})$ admits an orbifold pluri-harmonic metric $H_{\mathrm{orb},i}$. Since each $\mu_{\alpha}$ is étale over $W_{\mathrm{reg}}$, $H_{\mathrm{orb},i}$ descends to a pluri-harmonic metric $H_{i}$ on $(\mathcal{Q}_{W,i},\theta_{W,i})|_{W_{\mathrm{reg}}}$. By virtue of the uniqueness statement in Theorem 1.13 (1), the connection $D_{\theta_{W,i},H_{i}}$ coincides with the flat connection on $Q_{W,i}|_{W_{\mathrm{reg}}}$ and so $D_{\theta_{\mathrm{orb},i},H_{\mathrm{orb},i}}=D_{\mathrm{orb},i}.$
Note that the Hodge isomorphism theorem remains valid for compact complex orbifolds (cf. [78, Proposition 2.1]). Hence we may follow the argument of [23, Proof of Theorem A] to construct a Higgs orbi-bundle structure $(V_{\mathrm{orb}},\bar{\partial}_{\mathrm{orb}},\theta_{\mathrm{orb}})$ on $V_{\mathrm{orb}}$ such that each $V_{\mathrm{orb},i}$ is invariant under $\bar{\partial}_{\mathrm{orb}}$ and $\theta_{\mathrm{orb}}$, and the induced Higgs structure on each quotient $Q_{\mathrm{orb},i}$ coincides with $(Q_{\mathrm{orb},i},\bar{\partial}_{\mathrm{orb},i},\theta_{\mathrm{orb},i})$. Consequently,
@@
gives an extension of the collection of locally Higgs sheaves $\{\mu_{\alpha}^{*}(\mathcal{Q}_{W,i},\theta_{W,i})\}_{i=1}^{m}$. By applying [33, Lemma A.4] and noticing that $\mu_{\alpha}$ is étale over $W_{\mathrm{reg}}$, we see that $\left((\mu_{\alpha})_{*}(\mu_{\alpha}^{*}\mathcal{Q}_{W,i})\right)^{G_{\alpha}}$ is reflexive and coincides with the locally free sheaf $\mathcal{Q}_{W,i}$. Then $\mathcal{E}_{W}:=\left((\mu_{\alpha})_{*}(\mathcal{E}_{\mathrm{orb}})_{\alpha}\right)^{G_{\alpha}}$ is reflexive and admits a filtration by saturated subsheaves